# Diffusion Model-Based Image Editing: A Survey

### v2

@@ 7.1.2 Zero-Shot Learning
Context Prior Integration. This approach aims to extract inherent structures and textures from the uncorrupt part of an image as complementary to generate the content in the masked corrupt regions at each step, thereby maintaining global content consistency. Representatively, Repaint [266] builds on a pretrained unconditional diffusion model and modifies the standard denoising strategy by sampling the masked regions from the diffusion model and sampling the unmasked regions from the given image in each reverse step. GradPaint [267] also utilizes a pretrained diffusion model to estimate content for masked image regions. It then uses a custom loss function to measure coherence with unmasked parts, updating the generated content based on the loss’s gradient to ensure context consistency.
Degradation Decomposition. Image inpainting can be regarded as a specialized application of general linear inverse problems, which is represented as:
$\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n},$ (13)
@@
where $\mathbf{H}$ is the degradation operator (e.g., a mask matrix comprising 0 and 1 that indicates the corrupted regions of the image in inpainting), and $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$. The objective is to recover the original image $\mathbf{x}$ from the corrupted image $\mathbf{y}$. A few works implement it by performing decomposition on $\mathbf{H}$. DDRM [66], utilizing a pretrained denoising diffusion model and Singular Value Decomposition (SVD) of $\mathbf{H}$, transforms this recovery into an efficient iterative diffusion process within the spectral space of $\mathbf{H}$. Differently, DDNM [68] addresses the image inverse problem by applying a range-null space decomposition, where the image is divided into two distinct components via the pseudo-inverse of the degradation matrix $\mathbf{H}^{\dagger}$.
Posterior Edstimation. To solve the general noisy linear inverse problem as formulated in Eq. 13, several studies [70, 268, 269, 270, 271, 272] focus on estimating the posterior distribution $p(\mathbf{x}|\mathbf{y})$, leveraging the unconditional diffusion model based on Bayes’ theorem. It calculates the conditional posterior $p(\mathbf{x}_{t}|\mathbf{y})$ at each step of the reverse diffusion process. With Bayes’ theorem, it can be derived: $p(\mathbf{x}_{t}|\mathbf{y})=\frac{p(\mathbf{y}|\mathbf{x}_{t})p(\mathbf{x}_{t})}{p(\mathbf{y})}$. The corresponding gradient of the log-likelihood is then computed as:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}),$ (14)
where the score function $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ can be estimated using a pretrained score network $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$. The estimation of $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})$ then reduces to estimating the unknown term $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})$. Two representative earlier works, MCG [70] and DPS [268], approximate the posterior $p_{t}(\mathbf{y}|\mathbf{x}_{t})$ with $p_{t}(\mathbf{y}|\hat{\mathbf{x}}_{0})$, where $\hat{\mathbf{x}}_{0}$ is the noise-free expectation of $\mathbf{x}_{t}$, formalized as $\hat{\mathbf{x}}_{0}=E[\hat{\mathbf{x}}_{0}|\mathbf{x}_{t}]$ using Tweedie’s formula. Hence, Eq. 14 becomes:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})$ $=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$
$\approx-\rho\nabla_{\mathbf{x}_{t}}\|\mathbf{y}-\mathbf{H}\hat{\mathbf{x}}_{0}\|^{2}_{2}+\mathbf{s}_{\theta}(\mathbf{x}_{t},t),$ (15)
where $\rho=\frac{1}{\sigma^{2}}$ is the step size. MCG further adds a projection onto the measurement subspace post-update via Eq. 7.1.2, serving as a corrective measure for deviations from ideal data consistency. $\Pi$GDM [269] expands the posterior estimation framework in Eq. 14 to address linear, non-linear, and differentiable inverse problems by introducing the Moore-Penrose pseudoinverse $\mathbf{H}^{\dagger}$ of the degradation function $\mathbf{H}$. GDP [270] gives a different approximation formulated as:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})\approx-s\nabla_{\mathbf{x}_{t}}\mathcal{L}(\mathbf{H}\hat{\mathbf{x}}_{0}+\mathbf{n},\mathbf{y})-\lambda\nabla_{\mathbf{x}_{t}}\mathcal{Q}(\hat{\mathbf{x}}_{0}),$ (16)
where $\mathcal{L}$, $\mathcal{Q}$, $s$, and $\lambda$ are an image distance metric, a quality loss, a scale factor controlling the magnitude of guidance, and a scale factor for adjusting the quality of images, respectively. With a similar goal, CoPaint [271] uses the one-step estimation of the final generated image to predict $\hat{\mathbf{x}}_{0}$ deterministically to make the computation of the inpainting errors tractable at each intermediate denoising step.
@@ 7.2.1 Random Mask Training
Traditional image inpainting, which focuses on filling missing areas in images driven by surrounding context, often lacks precise control over the content. With the advance of text-to-image diffusion models, this limitation is being overcome by injecting additional user-provided multimodal conditions such as text descriptions, segmentation maps, and reference images as guidance. These models, e.g., Stable Diffusion and Imagen, can also be adapted to the inpainting task where the noisy background is replaced with a noisy version of the original image in the reverse diffusion process. This modification, however, sometimes results in unsatisfactory samples due to limited global context visibility during sampling. Addressing this, models like GLIDE [34] and Stable Inpainting (inpainting specialist v1.5 from Stable Diffusion [26]) further finetune the pretrained text-to-image models specialized for inpainting. They use a randomly generated mask along with the masked image and the full image caption, enabling the model to utilize information from the unmasked region during training. As pioneering works, they provide enlightening insights for subsequent research.
@@ 7.2.2 Precise Control Conditioning
An essential problem with random masking is that it may cover regions unrelated to the text prompt, leading the model to ignoring the prompt, especially when the masked regions are small or cover only a part of an object. To provide more precise control over the inpainted content, several approaches have been proposed. SmartBrush [150] introduces a precision factor to produce various mask types, from fine to coarse, by applying Gaussian blur to accurate instance masks. It allows users to choose between coarse masks, which include the desired object within the mask, and detailed masks that closely outline the object shape.
Imagen Editor [153] extends Imagen [27] through finetuning, employing a cascaded diffusion model for text-guided inpainting. It integrates image and mask context into each diffusion stage with three convolutional downsampling image encoders. The mask here is generated on-the-fly by an object detector, SSD Mobilenet v2 [273], instead of randomly masking.
PbE [142] and PhD [145] focus on using reference images for users’ customized inpainting. Specifically, PbE discards the text while only extracting subject information from the reference image using CLIP [242]. The reference image is cropped from the inpainting image itself with augmentations to support self-supervised training. In contrast, PhD integrates both text and reference image in the training pipeline same as PbE’s.
Uni-paint [155] first finetunes Stable Diffusion on the unmasked regions of the input images with null text and then uses various modes of guidance including unconditional, text-driven, stroke-driven, exemplar-driven inpainting, and a combination of them to perform multimodal inpainting during inference via different condition integration techniques. For facial inpainting, PVA [274] preserves identity by encoding a few reference images with the same identity as the masked image through Transformer blocks. The output image embedding, combined with the CLIP text embedding, then serves as extra conditions of the Stable Inpainting model through a parallel visual attention mechanism. SmartMask [154] utilizes semantic amodal segmentation data to create high-quality annotations for mask-free object insertion. It trains a diffusion model to predict instance maps from these annotations including both semantic maps and scene context. During inference, it uses panoptic segmentation to create a semantic map, which guides the prediction of object insertion masks.
@@ 7.2.3 Pretrained Diffusion Prior Exploiting
Blended Diffusion [229] employs CLIP to calculate the difference between the denoised image embedding from a pretrained unconditional diffusion model and the guided text embedding to refine the generation within masked regions. It then spatially blends the noisy masked regions with the corresponding noisy version of the unmasked part of the input image during each sampling step. Blended Latent Diffusion [222] takes this further by applying the technique in the LDM [26] for higher efficiency.
Inpaint Anything [231] combines SAM [246] and existing inpainting model such as LaMa [275] and Stable Inpainting [222] as a whole framework that allows users to remove, replace, or fill any part of an image with simple clicks, offering a high level of user-friendliness and flexibility in image editing. MagicRemover [186] starts with a null-text inversion [175] to project an image into a noisy latent code, and then optimizes subsequent codes during the sampling process. It introduces an attention-based guidance strategy and a classifier optimization algorithm to improve the precision and stability of the inpainting process with fewer steps.
@@ 8.1 Benchmark Construction
To address these problems, we introduce EditEval, a benchmark designed to evaluate general diffusion model-based image editing methods. EditEval includes a carefully curated dataset of 50 high-quality images, each accompanied by text prompts. EditEval evaluates performance on 7 common editing tasks selected from Table I. Additionally, we propose LMM Score, a quantitative evaluation metric that utilizes the capabilities of large multimodal models (LMMs) to assess the editing performance on different tasks. In addition to the objective evaluation provided by LMM Score, we also carry out a user study to incorporate subjective evaluation. The details of EditEval’s construction and its application are described as follows.
@@ 8.1.2 Dataset Construction
For image selection, we manually choose 50 images from Unsplash’s online repository of professional photographs, ensuring a wide diversity of subjects and scenes. These images are cropped to a square format, aligning with the input ratio for most editing models. We then categorize the chosen images into 7 groups, each associated with one of the specific editing tasks mentioned above.
@@ 8.2.1 Method Selection
For the evaluation of each editing task, we carefully select between 4 to 8 methods from Table I, encompassing a range of approaches including training-based, testing-time finetuning, and training & finetuning free methods. To ensure a fair and consistent comparison, our selection criteria are as follows: the method must require only text conditions, own the capability to handle specific tasks, and have open-source code available for implementation. We exclude domain-specific methods to avoid the constraints imposed by their limited applicability across diverse domains. These selected methods are specified in subsequent subsections.
@@ 8.2.2 Comparative Analysis
Performance Comparison. To present a thorough evaluation of the selected methods across the 7 selected editing tasks, we compute the mean and standard deviation of both $S_{LMM}$ and $S_{User}$ on all evaluation samples, as shown in Fig. 10. The color bars illustrate the range of scores from minimum to maximum. From the results, we can have several insights. First, there is no single method that outperforms others across all tasks. For example, while LEDITS++ [198] demonstrates commendable performance in object removal (Fig. 10(c)), it does not perform well in style change (Fig. 10(e)). Second, most methods exhibit a wide range of fluctuating scores in each task, as evidenced by the broad span between the minimum and maximum values and the large standard deviations of both $S_{LMM}$ and $S_{User}$. Such variability points to the unrobustness of these methods, suggesting that a method’s performance can be sample-dependent and may not consistently meet the editing requirements in various scenes. Third, some methods do exhibit stable and impressive performances for certain tasks. Null-Text Inversion (Null-Text) [175], for example, consistently shows high score distributions in object replacement, background change, and texture change, indicating its reliable [Uncaptioned image]
editing quality for some real-world applications. Last, the mean values of $S_{LMM}$ and $S_{User}$ are closely matched for each method within each task, suggesting that LMM Score well reflects users’ preferences. Although the methods’ performances fluctuate, the consistency between LMM Score and user perception across the tasks indicates that LMM Score is a reliable metric for image editing. Besides the quantitative results, we also present several visual examples for qualitative comparison in Fig. 13.
@@
Figure 12: Comparison of Pearson correlation coefficients between LMM Score / CLIPScore and the user study.
@@
LMM Score vs. CLIPScore. To further evaluate the effectiveness of LMM Score, we compute the Pearson correlation coefficients between a commonly used metric CLIPScore [279] and the user scores for comparison. CLIPScore is calculated using the specified text prompt $t_{edit}$ for a source sample $I_{src}$ and the corresponding edited image $I^{i}_{tgt}$ generated by method $i$ within a specific editing task [279]. Once all CLIPScore values are obtained for all samples across the 7 selected tasks, we calculate the Pearson correlation coefficients between these scores and the overall user scores on all the samples. Unlike LMM Score, CLIPScore does not consider specific factors and is thus directly compared with $S_{User}$. These coefficients are then averaged to yield an overall correlation coefficient for each task. Concurrently, we compute the Pearson correlation coefficients between the overall LMM scores $S_{LMM}$ and the overall user scores on all the samples and average them per task as well. All the $S_{LMM}$ scores are normalized to [0, 1] to match the range of CLIPScore. Fig. 12 visualizes these comparisons and demonstrates that for all these editing tasks, LMM Score maintains a closer correlation with user judgments than CLIPScore does, signifying a more accurate reflection of user evaluations and the effectiveness of LMM Score in the evaluation of various image editing tasks and methods.
@@ 9 Challenges and Future Directions
Despite the success achieved in image editing with diffusion models, there are still limitations that need to be addressed in future work.
Fewer-step Model Inference. Most diffusion-based models require a significant number of steps to obtain the final image during inference, which is both time-consuming and computationally costly, bringing challenges in model deployment and user experience. To improve the inference efficiency, few-step or one-step generation diffusion models have been studied [280, 281, 90]. Recent methods in this field reduce the number of steps by distilling knowledge from a pretrained strong diffusion model, so that the few-step model can mimic the behavior of the strong model. A more interesting yet challenging direction is to develop few-step models directly without relying on the pretrained models, such as consistency models [282, 283].
@@
Complex Object Structure Editing. Existing works can synthesize realistic colors, styles or textures when editing images. However, they still produce noticeable artifacts when dealing with complex structures, such as fingers, logos, and scene text. Attempts have been made to address these issues. Previous methods usually use negative prompting such as “six fingers, bad leg, etc.” to make the model avoid generating such images, which are effective in certain cases but not robust enough [287]. Recent works start to use layouts, edges or dense labels as guidance for editing the global or local structures of images [101, 103].
Complex Lighting and Shadow Editing. It is still challenging to edit the lighting or shadow of an object, which requires an accurate estimation of the lighting condition in the scene. Previous works such as Total Relighting [288] use a combination of networks to estimate the normal, albedo, and shading of the foreground object to obtain realistic relighting effects. Recently, diffusion-based models are proposed to edit the lighting of faces (DiFaReli [289]). However, the editing of portraits or general objects leveraging the strong lighting prior from the pretrained diffusion models remains an open area [290]. Similarly, ShadowDiffusion [62] explores diffusion-based shadow synthesis and can generate visually pleasing shadows of objects. However, accurately editing the shadow of an object under different background conditions using diffusion models is still an unsolved problem.

### v3

@@ 7.1.2 Zero-Shot Learning
Context Prior Integration. This type of methods extracts structures and textures from uncorrupted image parts to complement masked regions, ensuring global content consistency. Repaint [292] uses a pretrained unconditional diffusion model, modifying the denoising process by sampling masked areas from the model and unmasked areas from the image at each reverse step. Similarly, GradPaint [293] employs a pretrained diffusion model to estimate masked content, using a custom loss function to measure coherence with unmasked regions and updating the content based on the loss gradient.
@@
Degradation Decomposition. Image inpainting can be regarded as a specialized application of general linear inverse problems represented as $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$, where $\mathbf{H}$ is the degradation operator (e.g., a mask matrix indicating corrupted regions in inpainting), and $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$. The goal is to recover the original image $\mathbf{x}$ from the corrupted image $\mathbf{y}$. Some works focus on decomposition on $\mathbf{H}$. DDRM [53] uses a pretrained diffusion model and Singular Value Decomposition (SVD) of $\mathbf{H}$ to transform this recovery into an efficient iterative diffusion process in the spectral space of $\mathbf{H}$. Differently, DDNM [55] applys a range-null space decomposition, dividing the image into two distinct components via the pseudo-inverse of the degradation matrix.
Posterior Estimation. To solve the general noisy linear inverse problem, several studies [57, 294, 295, 296, 297, 298] focus on estimating the posterior distribution $p(\mathbf{x}|\mathbf{y})$, leveraging the unconditional diffusion model based on Bayes’ theorem. The conditional posterior $p(\mathbf{x}_{t}|\mathbf{y})$ is calculated at each step of the reverse diffusion process: $p(\mathbf{x}_{t}|\mathbf{y})=\frac{p(\mathbf{y}|\mathbf{x}_{t})p(\mathbf{x}_{t})}{p(\mathbf{y})}$. The corresponding gradient of the log-likelihood is then computed as $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$, where the score function $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ is estimated by a pretrained score network $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$. Thus, the problem reduces to estimating $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})$. DPS [294] approximates the posterior $p_{t}(\mathbf{y}|\mathbf{x}_{t})$ with $p_{t}(\mathbf{y}|\hat{\mathbf{x}}_{0})$. MCG [57] further introduces a projection onto the measurement subspace to ensure data consistency. $\Pi$GDM [295] extends this approach to both linear and non-linear inverse problems through the use of the Moore-Penrose pseudo-inverse $\mathbf{H}^{\dagger}$ of the degradation function $\mathbf{H}$. GDP [296] offers a different approximation strategy, incorporating an image distance metric and quality loss functions to guide the estimation process. Similarly, CoPaint [297] uses a one-step prediction of the final image to make inpainting errors more tractable during the denoising steps.
@@ 7.2.1 Random Mask Training
Traditional image inpainting, focusing on filling missing areas in images driven by surrounding context, often lacks precise control over the content. With the advance of text-to-image diffusion models, this limitation is being overcome by injecting additional multimodal conditions such as text descriptions, segmentation maps, and reference images. These models, e.g., Stable Diffusion and Imagen, can also be adapted to the inpainting task where the noisy background is replaced with a noisy version of the original image in the reverse diffusion process. However, this can sometimes result in unsatisfactory samples due to limited global context visibility. Addressing this, models like GLIDE [38] and Stable Inpainting (inpainting specialist v1.5 from Stable Diffusion [30]) further finetune the pretrained text-to-image models specialized for inpainting. They use a randomly generated mask along with the masked image and the full image caption, enabling the model to utilize information from the unmasked region during training. As pioneering works, they provide enlightening insights for subsequent research.
@@ 7.2.2 Precise Control Conditioning
An essential problem with random masking is that it may cover regions unrelated to the text prompt, leading the model to ignoring the prompt, especially when the masked regions are small. To provide more precise control, SmartBrush [173] introduces a precision factor to produce various mask types. It allows users to choose between coarse masks, which include the desired object within the mask, and detailed masks that closely outline the object shape. Imagen Editor [176] finetunes Imagen [31] by integrating image and mask context into each diffusion stage with three convolutional downsampling image encoders. The mask here is generated on-the-fly by an object detector, SSD Mobilenet v2 [299], instead of randomly masking.
PbE [165] and PhD [168] focus on using reference images for users’ personalized inpainting. Specifically, PbE uses CLIP to only extract subject information from the reference image, while PhD integrates both text and reference images in the training pipeline same as PbE’s. Uni-paint [178] first finetunes Stable Diffusion on the unmasked regions of the input images with null text and then uses various modes of guidance including unconditional, text-driven, stroke-driven, exemplar-driven inpainting, and a combination of them to perform multimodal inpainting during inference via different condition integration techniques. PVA [300] focuses on facial inpainting by encoding a few reference images with the same identity as the masked image through Transformer blocks. The encoded image embedding, combined with the CLIP text embedding, then serves as extra conditions of the Stable Inpainting model through a parallel visual attention mechanism. SmartMask [177] utilizes semantic amodal segmentation data to create high-quality annotations for mask-free object insertion. It trains a diffusion model to predict instance maps from these annotations including both semantic maps and scene context.
@@ 7.2.3 Pretrained Diffusion Prior Exploiting
Blended Diffusion [252] uses CLIP to compute the difference between the denoised image embedding from a pretrained diffusion model and the guided text embedding to refine generation within masked regions. It then blends the noisy masked regions with the noisy unmasked parts of the input image during each sampling step. Blended Latent Diffusion [245] takes this further by applying the technique in the LDM [30] for higher efficiency.
Inpaint Anything [254] combines SAM [272] with inpainting models like LaMa [301], providing a user-friendly framework for removing, replacing, or filling image parts with simple clicks. MagicRemover [209] employs null-text inversion [198] to project an image into a noisy latent code and then optimizes the codes during sampling. It introduces an attention-based guidance strategy and a classifier optimization algorithm to enhance inpainting precision and stability with fewer steps.
@@ 8.1 Benchmark Construction
To address these problems, we introduce EditEval, a benchmark designed to evaluate general diffusion model-based image editing methods. EditEval includes a carefully curated dataset of 150 high-quality images, each accompanied by text prompts. EditEval evaluates performance on 7 common editing tasks selected from Table I. Additionally, we propose LMM Score, a quantitative evaluation metric that utilizes the capabilities of large multimodal models (LMMs) to assess the editing performance on different tasks. In addition to the objective evaluation provided by LMM Score, we also carry out a user study to incorporate subjective evaluation. The details of EditEval’s construction and its application are described as follows.
@@ 8.1.2 Dataset Construction
For image selection, we manually choose 150 images from Unsplash’s online repository of professional photographs, ensuring a wide diversity of subjects and scenes. These images are cropped to a square format, aligning with the input ratio for most editing models. We then categorize the chosen images into 7 groups, each associated with one of the specific editing tasks mentioned above.
@@ 8.2.1 Method Selection
For the evaluation of each editing task, we carefully select between 4 to 8 methods from Table I, encompassing a range of [Uncaptioned image]
approaches including training-based, testing-time finetuning, and training & finetuning free methods. To ensure a fair and consistent comparison, our selection criteria are as follows: the method must require only text conditions, own the capability to handle specific tasks, and have open-source code available for implementation. We exclude domain-specific methods to avoid the constraints imposed by their limited applicability across diverse domains. These selected methods are specified in subsequent subsections.
@@ 8.2.2 Comparative Analysis
Performance Comparison. To present a thorough evaluation of the selected methods across the 7 selected editing tasks, we compute the mean and standard deviation of both $S_{LMM}$ and $S_{User}$ on all evaluation samples, as shown in Fig. 9. The color bars illustrate the range of scores from minimum to maximum. From the results, we can have several insights. First, there is no single method that outperforms others across all tasks. Second, most methods exhibit a wide range of fluctuating scores in each task, as evidenced by the broad span between the minimum and maximum values and the large standard deviations of both $S_{LMM}$ and $S_{User}$. Such variability points to the unrobustness of these methods, suggesting that a method’s performance can be sample-dependent and may not consistently meet the editing requirements in various scenes. Third, some methods do exhibit stable and impressive performances for certain tasks. Last, the mean values of $S_{LMM}$ and $S_{User}$ are closely matched for each method within each task, suggesting that LMM Score well reflects users’ preferences. Although the methods’ performances fluctuate, the consistency between LMM Score and user perception across the tasks indicates that LMM Score is a reliable metric for image editing. Besides the quantitative results, we also present several visual examples for qualitative comparison in Fig. 12.
@@
Figure 11: Comparison of Pearson correlation coefficients between four metrics and the user study.
@@
LMM Score vs. Other Metrics. To further assess the effectiveness of LMM Score, we compare it against three other metrics—CLIPScore [305], Directional CLIP Similarity [266], and TIFA Score [306]—by examining their Pearson correlation coefficients with the user study results. Unlike CLIPScore and Directional CLIP Similarity, which measure the feature distance between text and image directly in the CLIP feature space, TIFA Score evaluates the faithfulness of a generated image to its text input using visual question answering (VQA). Specifically, TIFA Score generates several question-answer pairs based on the text input using a language model and then determines image faithfulness by verifying whether the current VQA model can accurately answer these questions using the generated image. Fig. 11 presents these comparisons. It reveals that both CLIPScore and Directional CLIP Similarity show limited correlation with user judgments. While TIFA Score demonstrates better performance in semantic editing tasks such as object addition, it struggles with stylistic editing like style change. In contrast, LMM Score consistently exhibits the highest correlation with user evaluations across all tasks. This indicates that LMM Score more accurately captures the quality and relevance of image edits in alignment with human judgment, making it a more reliable metric for diverse editing scenarios.
@@ 9 Challenges and Future Directions
Despite the success achieved in image editing with diffusion models, there are still limitations to address in future work.
Fewer-Step Model Inference. Most diffusion-based models require a significant number of steps to obtain the final image during inference, which is both time-consuming and computationally costly, bringing challenges in model deployment and user experience. To improve the inference efficiency, few-step or one-step generation diffusion models have been studied [307, 308, 91]. Recent methods in this field reduce the number of steps by distilling knowledge from a pretrained strong diffusion model so that the few-step model can mimic the behavior of the strong model. A more interesting yet challenging direction is to develop few-step models directly without relying on the pretrained models, such as consistency models [309, 310].
@@
Complex Object Structure Editing. Existing works can synthesize realistic colors, styles or textures when editing images. However, they still produce noticeable artifacts when dealing with complex structures, such as fingers, logos, and scene text. Attempts have been made to address these issues. Previous methods usually use negative prompting such as “six fingers, bad leg, etc.” to make the model avoid generating such images, which are effective in certain cases but not robust enough [314]. Recent works start to use layouts, edges, or dense labels as guidance for editing the global or local structures of images [102, 104].
Complex Lighting and Shadow Editing. Editing the lighting or illumination of an image remains a challenging task [315], especially when aiming for realistic and consistent results. Traditional methods like Total Relighting [316] estimate normals, albedo, and shading to achieve realistic effects. Recent diffusion-based methods have advanced this area. For instance, DiffusionRig [317] and DiFaReli [318] estimate 3D parameters (e.g., shape, expression, SH lighting) from portraits and use these to condition diffusion models for face relighting. Other approaches [319, 320] reconstruct relightable 3D face models from a single image using diffusion models, enabling 3D editing. However, since these methods focus primarily on portrait relighting, their applicability to broader natural images is limited. More recent methods [321, 322, 323, 324] focus on training image-based, lighting-conditioned diffusion models on synthetic datasets. Specifically, Lasagna [321] is trained on a self-rendered 3D object dataset relit from multiple light source locations to learn two lighting layers for luminosity adjustment. Similarly, both DiLightNet [322] and Neural Gaffer [324] train light-conditioned ControlNet models on synthetic datasets, where 3D objects are rendered with different lighting using Blender. Although these methods are not restricted to portrait relighting, they still rely heavily on strong lighting priors from 3D models. LightIt [323] focuses on outdoor relighting using shading maps estimated by determining the Sun’s direction, but it is less suitable for indoor scenes. In contrast, Retinex-Diffusion [325] offers a training-free approach by reformulating the energy function of diffusion models based on the Retinex theory, allowing for lighting alterations without the need for 3D priors. Similarly, ShadowDiffusion [49] explores generating visually pleasing shadows of objects using degradation priors. However, accurately editing the lighting or shadow of an object under different background conditions using diffusion models remains an unsolved problem.
@@
High-Resolution Image Generation and Editing. High-resolution images, typically those with resolutions of 1024x1024 pixels or higher, have become increasingly important in various applications. While several diffusion-based models focus on high-resolution image generation, such as SDXL [93] and PixArt [94], specialized diffusion models for high-resolution image editing are relatively scarce. Some studies [328, 213, 210] have claimed the capability to perform high-resolution image editing, but few have demonstrated concrete results. One of the key challenges in high-resolution image editing is ensuring precise modifications. The model must achieve a high degree of accuracy, particularly in tasks involving pixel-level adjustments and edge refinement, where maintaining fine details is essential. This level of precision becomes increasingly demanding with higher resolutions. Another major challenge is maintaining computational efficiency. As resolution increases, the demand for GPU memory and processing power grows substantially, which puts pressure on both hardware resources and the model’s runtime performance. Therefore, it becomes essential to strike a balance between precision and speed to ensure the editing process remains efficient and smooth.